# Domain-aware priors enable vertical federated learning in data-scarce coral multi-omics Sam Victor1*

### v1

@@ Abstract
Vertical federated learning (VFL) enables multi-laboratory collaboration on distributed multi-omics datasets without sharing raw data, but exhibits severe instability under extreme data scarcity () when applied generically. Here, we investigate how domain-aware design choices; specifically gradient saliency–guided feature selection with biologically motivated priors; affect the stability, interpretability, and failure modes of VFL architectures in small-sample coral stress classification (N = 13 samples, P = 90,579 features across transcriptomics, proteomics, metabolomics, and microbiome data). We benchmark a domain-aware VFL framework against two baselines on the Montipora capitata thermal stress dataset: (i) a standard NVFlare-based VFL and (ii) LASER, a state-of-the-art label-aware VFL method. Domain-aware VFL achieves an AUROC of 0.833 ± 0.030 after reducing dimensionality by 98.6% (90,579 to 1,300 features), significantly outperforming NVFlare VFL, which performs at chance level (AUROC 0.500 ± 0.125, p = 0.0058, Cohen’s d = 2.684). While LASER shows modest improvement (AUROC 0.600 ± 0.215), it exhibits substantially higher variance and does not reach statistical significance (p = 0.095) despite a large effect size. We show that domain-aware feature selection yields highly stable top-feature sets across multiple analysis parameters, supporting reproducible and interpretable model behavior. Negative control experiments using permuted labels produce AUROC values below chance (0.262), confirming the absence of data leakage and indicating that observed performance arises from genuine biological signal. Together, these results motivate empirically grounded design principles for VFL in extreme regimes, emphasizing the importance of domain-informed dimensionality reduction, stability-focused evaluation, and interpretable feature selection for learning from scarce biological data.
@@
Coral reefs are increasingly threatened by climate-driven heat stress, but understanding how corals respond at the molecular level requires integrating many different types of biological data, including genes, proteins, metabolites, and associated microbes. These measurements are often generated by different laboratories and are difficult to analyze together, especially when only a small number of coral samples are available.
In this study, we address this challenge by developing a privacy-preserving federated learning framework that allows multiple research groups to collaboratively analyze multi-omic coral data without sharing their raw datasets. To make learning feasible in this data-scarce setting, we incorporate existing biological knowledge about coral stress responses to guide the selection of informative molecular features.
@@ Introduction
Coral reefs face unprecedented collapse from climate-driven thermal stress, with mass bleaching events causing global reefs to experience mortality since 2016. [1]. Understanding molecular stress responses at the holobiont level (integrating coral host, symbiotic dinoflagellates, and microbiome) requires multi-omics profiling across transcriptomics, proteomics, metabolomics, and microbial community composition [3]. However, the field confronts a fundamental constraint: coral sampling is ecologically restricted (N=13 samples typical), while modern omics platforms generate extreme feature dimensionality (P=90,579 features), creating a catastrophic PN regime (P/N ratio 6,000) where traditional machine learning fails completely [4]. Simultaneously, institutional barriers prevent data pooling (laboratories protect proprietary datasets due to intellectual property concerns, publication precedence, and data sovereignty policies) necessitating privacy-preserving collaborative approaches [3].
@@
The challenge of analyzing distributed, sensitive biological data has led to the emergence of privacy-preserving machine learning (PPML) as a core necessity in bioinformatics [6]. In the context of multi-omics, specific frameworks such as PPML Omics have been proposed to facilitate secure collaboration across institutions while maintaining data confidentiality [7]. The application of federated learning to transcriptomic and proteomic data is increasingly common, though often limited by the synchronization costs associated with heterogeneous datasets [8]. Furthermore, the integration of microbial community data into these models is supported by evidence that the microbiome plays a critical role in host thermal tolerance [9]. Vertical federated learning (VFL) specifically addresses cases where features are split across different holders, a scenario frequently encountered in clinical and ecological multi-omics [10]. Despite the architectural advantages of VFL, existing literature highlights that its performance is sensitive to the alignment of latent representations across silos [11]. This instability is particularly pronounced when dealing with the vast feature spaces of modern sequencing technologies [12]. Recent studies have emphasized that the inclusion of biological priors can serve as a regularizer in high-dimensional genomic modeling [13]. The shift toward domain-aware architectures is also motivated by the need for regulatory compliance in global research networks [14]. Ultimately, the development of stable VFL frameworks for small-N cohorts remains an open area of research necessary for precision conservation [15].
@@
Multi-omics integration offers transformative potential for coral conservation: by jointly analyzing transcriptomic regulation, proteomic effector responses, metabolomic phenotypes, and microbiome restructuring, researchers can identify diagnostic biomarkers of bleaching susceptibility, predict mortality risk before visible symptoms, and discover intervention targets [16]. Machine learning provides the only viable framework for extracting patterns from this high-dimensional, multi-modal data. No human expert can manually integrate 90,579 molecular features across four biological scales. Recent successes include neural network-based prediction of coral thermotolerance from gene expression and ensemble models identifying metabolic markers of symbiont breakdown. Yet this promise remains unrealized at scale due to the extreme PN problem. The Montipora capitata thermal stress dataset exemplifies the challenge: 13 biological replicates measured across 90,579 features (62,038 transcripts, 4,054 proteins, 12,055 metabolites, 12,432 microbial OTUs). Standard machine learning algorithms (support vector machines, random forests, deep neural networks) require NP or at minimum NP for stable parameter estimation. When P/N1, models overfit catastrophically, memorizing training noise rather than learning generalizable stress signatures. Regularization (L1/L2 penalties, dropout) provides marginal relief but cannot overcome thousandfold parameter-to-sample ratios.
@@
Vertical Federated Learning (VFL) enables multiple institutions holding different feature sets (omics layers) for overlapping samples to collaboratively train models without sharing raw data. Each laboratory (silo) encodes its local omics layer into privacy-preserving embeddings transmitted to a central server for fusion, while gradients flow back for local encoder updates. This architecture perfectly suits coral research: transcriptomics at genomics facilities, proteomics at mass spectrometry cores, metabolomics at specialized labs, microbiome sequencing at separate institutions. All contributing to a unified stress classifier without revealing proprietary measurements. However, existing VFL methods exhibit observed failure patterns commonly reported in PN settings:
@@
State-of-the-art methods like LASER-VFL attempt to learn label-aware aligned representations across modalities. With abundant data (N10,000), LASER achieves 1-3% AUROC improvements over baselines. But in PN regimes, its latent space alignment tries to align noise with noise. Without prior dimensionality reduction, the model cannot distinguish signal-bearing from noise-bearing features, causing training instability (high variance across random seeds) and modest performance gains (AUROC0.6).
@@
We hypothesize that domain-aware design choices (specifically, gradient saliency-guided feature selection with biological layer weighting) can mitigate VFL failure patterns under PN constraints. The key insight: decades of coral stress physiology provide strong priors. Not all 90,579 features contribute equally. Specific pathways (heat shock response, oxidative stress enzymes, photosystem damage markers) are mechanistically implicated in bleaching, while most genomic/metabolomic background remains unchanged under stress. Gradient-based saliency methods using neural network encoder backpropagation quantify per-feature importance: which molecular measurements most influence stress classification? By applying biological layer weights (1.5× for transcriptomics as hypothesized primary regulatory driver, 1.0× proteomics baseline, 0.8× metabolomics, 0.5× microbiome) before ranking and selecting top-K features, we impose domain structure that generic VFL ignores. This reduces dimensionality from 90,579 to 1,300 features (98.6%) before federated training, potentially:
1.
Restoring favorable signal-to-noise ratios: P/N drops from 6,967 to 100, enabling gradient-based optimization to identify stable decision boundaries.
2.
Stabilizing training dynamics: Biologically-grounded feature selection may produce consistent feature subsets across random seeds, reducing variance.
3.
Enabling interpretability: Selected features map to known stress pathways, validating model predictions against mechanistic biology.
@@
Failure mode characterization: Quantitative demonstration that NVFlare VFL exhibits gradient noise domination (AUROC 0.500±0.125, near-constant predictions) and LASER-VFL suffers representation collapse (AUROC 0.600±0.215, 7× higher variance than domain-aware approach).
@@
Stability as primary outcome: Domain-aware VFL achieves 7-fold variance reduction relative to LASER (SD 0.030 vs 0.215) and 4-fold reduction relative to NVFlare (SD 0.030 vs 0.125), with perfect feature ranking stability (Jaccard=1.0 across ), establishing that stability, not peak performance, is the critical metric in small-sample regimes.
@@
Validation rigor: Negative control experiments with permuted labels yield AUROC 0.262 (anti-learning phenomenon), confirming genuine biological signal and absence of data leakage, with true-to-random performance ratio (0.833/0.262=3.18) substantially exceeding random guessing baseline.
@@
Design principles for PN VFL: Three empirically motivated insights: (1) biological priors for dimensionality reduction appear necessary in this regime; (2) stability metrics matter more than peak performance when N100; (3) interpretable feature selection mitigates gradient noise domination in our setting.
This work provides the first systematic analysis of VFL failure modes in extreme PN regimes and demonstrates that domain-aware architectures enable privacy-preserving multi-laboratory collaboration on scarce coral omics data, a critical capability for global conservation research.
@@
We utilized the publicly available Montipora capitata (rice coral) thermal stress multi-omics dataset from Stephens et al. (DOI: 10.5281/zenodo.6861688). This dataset comprehensively characterizes coral holobiont responses to experimental heat stress through four integrated omics layers.
@@
Data S2. Processed metabolite data from the four M. capitata colonies (n=3 per treatment/time point/colony). Accumulation values (unnormalized) and peak counts, and compounds IDs are shown for each metabolite identified.
Data S3. Processed meta-transcriptome data for M. capitata colony MC289 (n=3 per treatment/time point). Read counts (unnormalized) for each of the predicted proteins in the M. capitata genome is shown for each of the samples.
Data S4. Processed microbiome 16S rRNA data from the four M. capitata colonies (n=3 per treatment/time point/colony). The (unnormalized) number of reads assigned to each ASV are shown for each of the samples.
Data S5. The (unnormalized) number of reads assigned to each of the ASV that survived filtering are shown for each of the samples.
@@
Preprocessing pipeline:
Raw data underwent standardized preprocessing: (1) Excel sheets extracted to tab-separated values (TSV) using automated scripts preserving numerical precision, (2) feature names standardized to remove special characters while retaining biological identifiers, (3) sample alignment performed by computing the intersection of available samples across all four omics layers, yielding 13 common samples (21 transcriptomics, 14 proteomics, 84 metabolomics, 83 microbiome reduced to 13 shared specimens), (4) microbiome features filtered to retain only OTUs with 0.1% relative abundance, reducing dimensionality from 27,807 to 12,432 features, and (5) StandardScaler normalization (zero mean, unit variance) applied per-omics during training, with scalers fitted exclusively on training data to prevent information leakage. Missing values (131 in proteomics) were replaced with zero. No log transformation or additional feature selection occurred during preprocessing to preserve raw abundance information and ensure fair comparison across methods.
@@
Encoder architecture. Each omics encoder comprises: (1) input layer matching feature dimensionality (), (2) progressive dimension reduction through hidden layers (input 128 64), where is computed as the geometric mean between input and 128 dimensions, (3) layer normalization after each linear transformation, (4) ReLU activation and dropout (p=0.3) for regularization, and (5) final 64-dimensional embedding output. Fusion head contains: (1) concatenated embedding input (256 dimensions = 4 silos 64), (2) two hidden layers (256 128 64) with layer normalization, ReLU activation, and dropout (p=0.2), and (3) output layer with sigmoid activation for binary classification.
@@
Domain-aware VFL with gradient saliency-guided feature selection
The critical innovation enabling domain-aware VFL for small-sample data is neural network gradient-based feature pruning prior to federated training.
Gradient saliency computation via encoder backpropagation. For each omics layer independently: (1) initialize a shallow neural network encoder (input 128 64 dimensions) matching the VFL architecture, (2) perform forward pass on training samples through the encoder to generate embeddings, (3) compute gradients via backpropagation by differentiating each embedding dimension with respect to input features, (4) accumulate absolute gradients across all embedding dimensions to obtain per-feature attribution scores, and (5) aggregate per-feature importance as mean absolute gradient: , where is the encoder and is feature for sample . This encoder-intrinsic approach computes feature importance directly from the neural network that will be used in VFL training, avoiding surrogate model approximations and ensuring attribution consistency with the federated architecture.
Biological weighting scheme. To test whether upstream regulatory mechanisms (gene expression) drive thermal stress classification and to yield a more domain-aware model, we applied differential weights to raw gradient saliency scores before feature ranking. The weighting scheme reflected established biological priors about information flow in stress response cascades:
@@
This a priori weighting favored transcriptomics as the hypothesized primary driver, with downstream layers progressively down-weighted based on their distance from regulatory control.
Top-K feature selection. For each omics layer: (1) multiply raw saliency scores by layer-specific weights, (2) rank features by descending weighted importance, (3) select top-K features with highest weighted saliency scores, and (4) filter both training and test data to retain only selected features. We tested K and selected K=500 based on performance-complexity tradeoff. This reduces total dimensionality from 90,579 to 1,300 features. The selected subset retains essential stress-discriminative information while eliminating noise that exacerbates overfitting in small-sample scenarios.
@@
Domain-aware VFL training. After feature selection: (1) split training into warmup phase (20 rounds, standard VFL on selected features to initialize embeddings) and refinement phase (30 rounds, continued training with lower learning rate), (2) train encoders on reduced feature sets significantly improving gradient stability and convergence, and (3) maintain identical communication protocol to standard VFL for fair comparison.
@@
We compared domain-aware VFL against two baselines:
NVFlare VFL (Base VFL). Standard NVIDIA FLARE vertical federated learning using the authentic NVFlare 2.4 API with DXO (Data eXchange Object) and Shareable communication patterns. This represents industry-standard VFL without domain-specific modifications. Trained for 50 federated rounds with 1 local epoch per round on full 90,579 features.
@@
Compute budget matching. All methods used matched optimizer step budgets to ensure fair comparison: NVFlare VFL (50 rounds × 1 local epoch = 50 steps), Domain-Aware VFL (50 rounds total = 50 steps), LASER (50 epochs = 50 steps). Budget ratio VFL/LASER = 1.00 (within target range 0.9-1.1).
Hyperparameter parity. Equivalent hyperparameters across methods: batch size (8), hidden dimensions (64-128), learning rates (5e-5 for VFL clients, 5e-4 for VFL server, 1e-3 for LASER), weight decay (1e-4), early stopping patience (15 rounds/epochs).
@@
Statistical testing. Paired t-tests assessed significance between methods using matched random seeds. Cohen’s d quantified effect sizes (negligible 0.2, small 0.2-0.5, medium 0.5-0.8, large 0.8). Wilcoxon signed-rank test provided non-parametric validation.
Negative control experiments. To validate absence of data leakage and confirm genuine biological signal, we performed negative control experiments where stress labels were randomly permuted before training. Domain-aware VFL was trained on shuffled labels using identical protocols (feature selection, cross-validation, hyperparameters) as true-label experiments. Performance on permuted labels provides a baseline for comparison against true-label results.
Saliency ranking stability analysis. Feature ranking stability evaluated by computing Jaccard similarity between top-20 features selected at different K thresholds (200, 500, 1000). Perfect stability (Jaccard=1.0) indicates robust domain knowledge extraction.
@@
All code implemented in Python 3.10 using PyTorch 2.9.1 [20] (neural networks and gradient-based feature importance), NVIDIA FLARE 2.4 [17] (VFL infrastructure), scikit-learn 1.7.2 [21] (preprocessing, metrics), NumPy 1.23, and pandas 1.5. Experiments were conducted on consumer-grade hardware (11th Gen Intel Core i7-11800H @ 4.60 GHz, 16 GB RAM, Debian Linux 13, kernel 6.12.57) to demonstrate computational accessibility without specialized GPU requirements. Complete codebase, preprocessed data, and configuration files available at ”https://github.com/samvictordr/domain-aware-vfl”.

### v3

# Domain-aware priors stabilize, not merely enable, vertical federated learning in data-scarce coral multi-omics Sam Victor1*

@@ Abstract
Vertical federated learning (VFL) enables multi-laboratory collaboration on distributed multi-omics datasets without sharing raw data, but exhibits severe instability under extreme data scarcity () when applied generically. Here, we investigate how domain-aware design choices—specifically gradient saliency-guided feature selection with biologically motivated priors—affect the stability, interpretability, and failure modes of VFL architectures in small-sample coral stress classification (N = 13 samples, P = 90,579 features across transcriptomics, proteomics, metabolomics, and microbiome data). We benchmark REEF, a domain-aware VFL framework, against two baselines on the Montipora capitata thermal stress dataset: (i) a standard NVFlare-based VFL and (ii) LASER, a state-of-the-art label-aware VFL method. REEF achieves an AUROC of 0.776 ± 0.039 after reducing dimensionality by 98.6% (90,579 to 1,300 features), substantially outperforming NVFlare VFL at chance level (AUROC 0.500 ± 0.125, , Cohen’s ) and numerically exceeding LASER (AUROC 0.557 ± 0.191, , Cohen’s ), with 3–5-fold variance reduction (SD 0.039 vs 0.125 and 0.191). An equal-weights ablation holding feature count constant at 1,300 but removing domain priors confirms that biological priors specifically contribute stability: removing priors yields statistically indistinguishable mean AUROC () but 2.3 higher variance (CV 0.110 vs 0.050). Domain-aware feature selection yields deterministic, budget-invariant feature rankings (Jaccard=1.0 across selection thresholds). Negative control experiments using permuted labels produce AUROC near or below chance (0.357 for REEF, 0.238 for NVFlare VFL), consistent with the absence of gross data leakage; however, with , this control rules out inflated performance rather than formally confirming below-chance behavior. These results motivate design principles for VFL in extreme regimes, emphasizing domain-informed dimensionality reduction, stability-focused evaluation, and interpretable feature selection for scarce biological data.
@@
Coral reefs are increasingly threatened by climate-driven heat stress, but understanding how corals respond at the molecular level requires integrating multiple types of biological data, such as genes, proteins, metabolites, and microbes. These measurements are often generated by separate laboratories and difficult to combine, especially when only a small number of coral samples are available.
In this study, we address this challenge by developing a privacy-preserving federated learning framework enabling multiple research groups to collaboratively analyze multi-omic coral data without sharing raw data. To address data scarcity, we incorporate biological knowledge about coral stress responses to guide the selection of informative molecular features.
@@ Introduction
Coral reefs face unprecedented collapse from climate-driven thermal stress, with mass bleaching events causing global reefs to experience mortality since 2016 [1]. Understanding molecular stress responses at the holobiont level (integrating coral host, symbiotic dinoflagellates, and microbiome) requires multi-omics profiling across transcriptomics, proteomics, metabolomics, and microbial community composition [3]. However, the field confronts a fundamental constraint: coral sampling is ecologically restricted (N=13 samples typical), while modern omics platforms generate extreme feature dimensionality (P=90,579 features), creating a severe PN regime (P/N ratio 6,000) where traditional machine learning fails completely [4]. Simultaneously, institutional barriers prevent data pooling (laboratories protect proprietary datasets due to intellectual property concerns, publication precedence, and data sovereignty policies) necessitating privacy-preserving collaborative approaches [3].
@@
The challenge of analyzing distributed, sensitive biological data has led to the emergence of privacy-preserving machine learning (PPML) as a core necessity in bioinformatics. In the context of multi-omics, specific frameworks such as PPML Omics have been proposed to facilitate secure collaboration across institutions while maintaining data confidentiality [6, 7]. Centralized graph-based approaches such as MOGONET [5] have demonstrated the potential of multi-omics integration for biomedical classification, but these methods require pooling raw data and thus cannot satisfy institutional privacy constraints. Furthermore, the integration of microbial community data into these models is supported by evidence that the microbiome plays a critical role in host thermal tolerance [8]. Vertical federated learning (VFL) specifically addresses cases where features are split across different holders, a scenario frequently encountered in clinical and ecological multi-omics [9]. Despite the architectural advantages of VFL, existing literature highlights that its performance is sensitive to the alignment of latent representations across silos [10]. Recent studies have emphasized that the inclusion of biological priors can serve as a regularizer in high-dimensional genomic modeling [11]. Ultimately, the development of stable VFL frameworks for small-N cohorts remains an open area of research necessary for precision conservation [13].
@@
Multi-omics integration offers transformative potential for coral conservation: by jointly analyzing transcriptomic regulation, proteomic effector responses, metabolomic phenotypes, and microbiome restructuring, researchers can identify diagnostic biomarkers of bleaching susceptibility, predict mortality risk before visible symptoms, and discover intervention targets [14]. The Montipora capitata thermal stress dataset exemplifies the challenge: 13 biological replicates measured across 90,579 features (62,038 transcripts, 4,054 proteins, 12,055 metabolites, 12,432 microbial OTUs). Standard machine learning algorithms (support vector machines, random forests, deep neural networks) require NP or at minimum NP for stable parameter estimation. When P/N1, models overfit severely, memorizing training noise rather than learning generalizable stress signatures.
@@
Vertical Federated Learning (VFL) enables multiple institutions holding different feature sets (omics layers) for overlapping samples to collaboratively train models without sharing raw data. Each laboratory (silo) encodes its local omics layer into privacy-preserving embeddings transmitted to a central server for fusion, while gradients flow back for local encoder updates. This architecture perfectly suits coral research: transcriptomics at genomics facilities, proteomics at mass spectrometry cores, metabolomics at specialized labs, microbiome sequencing at separate institutions—all contributing to a unified stress classifier without revealing proprietary measurements. However, existing VFL methods exhibit observed failure patterns commonly reported in PN settings:
@@
State-of-the-art methods like LASER-VFL attempt to learn label-aware aligned representations across modalities. With abundant data (N10,000), LASER achieves 1-3% AUROC improvements over baselines. But in PN regimes, its latent space alignment tries to align noise with noise. Without prior dimensionality reduction, the model cannot distinguish signal-bearing from noise-bearing features, causing training instability (high variance across random seeds) and modest performance gains (AUROC0.56).
@@
We hypothesize that domain-aware design choices (specifically, gradient saliency-guided feature selection with biological layer weighting) can mitigate VFL failure patterns under PN constraints. The key insight: decades of coral stress physiology provide strong priors. Not all 90,579 features contribute equally. Specific pathways (heat shock response, oxidative stress enzymes, photosystem damage markers) are mechanistically implicated in bleaching, while most genomic/metabolomic background remains unchanged under stress. Gradient-based saliency methods using neural network encoder backpropagation quantify per-feature importance: which molecular measurements most influence stress classification? We impose domain structure that generic VFL ignores through two mechanisms: omic-layer feature budgets (controlling how many features are selected per layer) and biological weights applied to encoder embeddings during training (1.5× transcriptomics, 1.0× proteomics, 0.8× metabolomics, 0.5× microbiome), reflecting hypothesized regulatory importance in the stress response cascade. This reduces dimensionality from 90,579 to 1,300 features (98.6%) before federated training, with three hypothesized benefits: restoring a tractable P/N ratio (6,967 → 100), stabilizing gradient updates by selecting biologically consistent feature subsets, and enabling interpretability by retaining features with known mechanistic roles in coral stress pathways.
@@
Failure mode characterization: Quantitative demonstration that NVFlare VFL exhibits gradient noise domination (AUROC 0.500±0.125, near-constant predictions) and LASER-VFL suffers representation collapse (AUROC 0.557±0.191, approximately 4.9× higher variance than domain-aware approach).
@@
Stability as primary outcome: REEF achieves 3–5-fold variance reduction relative to both LASER (SD 0.039 vs 0.191) and NVFlare (SD 0.039 vs 0.125), with deterministic, budget-invariant feature ranking (Jaccard=1.0 across ), a property that follows from computing importance scores once on a fixed training set (see Methods). An equal-weights ablation holding feature count constant at 1,300 confirms that biological priors specifically drive the stability advantage: removing priors while retaining aggressive dimensionality reduction yields statistically indistinguishable mean AUROC () but 2.3 higher variance (CV 0.110 vs 0.050), establishing that stability, not peak performance, is the critical metric in small-sample regimes and that domain knowledge is its source.
@@
Validation rigor: Negative control experiments with permuted labels yield AUROC 0.357 (domain-aware) and 0.238 (NVFlare), both near or below chance, providing evidence against gross data leakage, with true-to-random performance ratio (0.776/0.357=2.17); however, with only samples, sub-chance permuted-label AUROC values remain within normal sampling variation, so this control is necessary but not sufficient to rule out all overfitting.
@@
Design principles for PN VFL: Three empirically motivated insights: (1) aggressive dimensionality reduction enables above-chance learning in this regime, while biological priors are specifically necessary for stable learning, as confirmed by ablation; (2) stability metrics matter more than peak performance when N100; (3) interpretable feature selection mitigates gradient noise domination in our setting.
This work provides the first systematic analysis of VFL failure modes in extreme PN regimes, demonstrating that privacy-preserving multi-laboratory collaboration on scarce coral omics data is achievable when dimensionality reduction restores statistical tractability and domain knowledge stabilizes the learned representations.
@@
We utilized the publicly available Montipora capitata (rice coral) thermal stress multi-omics dataset from Stephens et al. (DOI: 10.5281/zenodo.6861688) [3, 2]. This dataset comprehensively characterizes coral holobiont responses to experimental heat stress through four integrated omics layers.
@@
Data S2. Processed metabolite data from the four M. capitata colonies (n=3 per treatment/time point/colony). Accumulation values (unnormalized) and peak counts, and compound IDs are shown for each metabolite identified.
Data S3. Processed meta-transcriptome data for M. capitata colony MC289 (n=3 per treatment/time point). Read counts (unnormalized) for each of the predicted proteins in the M. capitata genome are shown for each of the samples.
Data S4. Processed microbiome 16S rRNA data from the four M. capitata colonies (n=3 per treatment/time point/colony). The (unnormalized) number of reads assigned to each ASV is shown for each of the samples.
Data S5. The (unnormalized) number of reads assigned to each of the ASVs that survived filtering is shown for each of the samples.
@@
Preprocessing pipeline.
Raw data underwent standardized preprocessing: (1) Excel sheets extracted to tab-separated values (TSV) using automated scripts preserving numerical precision, (2) feature names standardized to remove special characters while retaining biological identifiers, (3) sample alignment performed by computing the intersection of available samples across all four omics layers, yielding 13 common samples (21 transcriptomics, 14 proteomics, 84 metabolomics, 83 microbiome reduced to 13 shared specimens), (4) microbiome features filtered to retain only OTUs with 0.1% relative abundance, reducing dimensionality from 27,807 to 12,432 features, and (5) StandardScaler normalization (zero mean, unit variance) applied per-omics during training, with scalers fitted exclusively on training data to prevent information leakage. Missing values (123 in proteomics) were replaced with zero. No log transformation or additional feature selection occurred during preprocessing to preserve raw abundance information and ensure fair comparison across methods.
@@
Encoder architecture. Each omics encoder comprises: (1) input layer matching feature dimensionality (), (2) progressive dimension reduction through hidden layers (input 128 64), where provides progressive dimension reduction (yielding for high-dimensional omics and for low-dimensional layers), (3) layer normalization after each linear transformation, (4) ReLU activation and dropout (p=0.3) for regularization, and (5) final 64-dimensional embedding output. Fusion head contains: (1) concatenated embedding input (256 dimensions = 4 silos 64), (2) two hidden layers (256 128 64) with layer normalization, ReLU activation, and dropout (p=0.3), and (3) output layer with sigmoid activation for binary classification.
@@
REEF: gradient saliency-guided feature selection for robust federated learning
The critical innovation enabling REEF (Robust Expert Encoder Federation) for small-sample data is neural network gradient-based feature pruning prior to federated training.
Gradient saliency computation via encoder backpropagation. For each omics layer independently: (1) initialize the full omics encoder (input 128 64 dimensions, as described above) with Xavier-normal weights (gain=0.1); (2) execute a supervised warmup phase of 20 VFL training rounds in which encoder weights are updated via backpropagation through the full VFL fusion head using binary cross-entropy loss against thermal stress labels, shaping the encoder’s activation landscape toward stress-discriminative representations; (3) compute a single Jacobian saliency pass on the warmup-trained encoder: for each of the 64 embedding dimensions , compute by calling .backward() on across all training samples; (4) accumulate absolute gradients across all embedding dimensions to obtain per-feature attribution scores; and (5) aggregate per-feature importance as mean absolute gradient: , where is the warmup-trained encoder and is feature for sample . This warmup-informed, encoder-intrinsic approach ensures that saliency scores reflect gradient dynamics shaped by supervised stress classification, while the Jacobian computation itself remains label-free and architecture-consistent with the subsequent federated training phase, avoiding surrogate model approximations.
Biological weighting scheme. To incorporate biological priors about information flow in stress response cascades, we implemented domain-aware weighting at two levels. At the feature selection level, per-omic topk ratios control the feature budget allocated to each omic layer (transcriptomics , proteomics , metabolomics , microbiome ). At the training level, encoder embeddings are scaled by biological weights reflecting hypothesized regulatory importance as set in Table 2.
@@
These embedding weights amplify or attenuate each omic layer’s contribution during the forward pass, with transcriptomics favored as the hypothesized primary driver and downstream layers progressively down-weighted.
Top-K feature selection. For each omics layer: (1) compute raw gradient saliency scores via encoder backpropagation, (2) rank features by descending importance, (3) select the top features per omic layer using the layer-specific topk ratio , and (4) filter both training and test data to retain only selected features. We tested K and selected K=500 based on performance-complexity tradeoff. With K=500, per-omic allocations are: transcriptomics 150, proteomics 250, metabolomics 500, microbiome 400, totaling 1,300 features. This reduces total dimensionality from 90,579 to 1,300 features (98.6% reduction). The selected subset retains essential stress-discriminative information while eliminating noise that exacerbates overfitting in small-sample scenarios.
@@
REEF training. After feature selection: (1) split training into warmup phase (20 rounds, standard VFL on selected features to initialize embeddings) and refinement phase (30 rounds, continued training with selected features), (2) train encoders on reduced feature sets significantly improving gradient stability and convergence, and (3) maintain identical communication protocol to standard VFL for fair comparison.
@@
We compared REEF against two baselines:
NVFlare VFL (Base VFL). Standard NVIDIA FLARE vertical federated learning using the authentic NVFlare 2.7.1 API with DXO (Data eXchange Object) and Shareable communication patterns. This represents industry-standard VFL without domain-specific modifications. Trained for 50 federated rounds with 1 local epoch per round on full 90,579 features.
@@
Compute budget matching. All methods used matched optimizer step budgets to ensure fair comparison: NVFlare VFL (50 rounds × 1 local epoch = 50 steps), REEF (50 rounds total = 50 steps), LASER (50 epochs = 50 steps). Budget ratio VFL/LASER = 1.00 (within target range 0.9-1.1).
Hyperparameter parity. Equivalent hyperparameters across methods: batch size (8), hidden dimensions (64-128), learning rates (5e-5 for VFL clients, 5e-4 for VFL server, 1e-3 for LASER), weight decay (1e-4 for VFL, 1e-5 for LASER), early stopping patience (15 rounds/epochs).
@@
Statistical testing. Paired -tests assessed significance between methods using matched random seeds ( seed pairs). Cohen’s was computed as the mean of per-seed differences divided by their population standard deviation (, treating the five seeds as the complete experiment rather than a sample from a larger population); thresholds follow standard conventions (negligible 0.2, small 0.2–0.5, medium 0.5–0.8, large 0.8). Wilcoxon signed-rank tests provided non-parametric validation, although with paired observations the minimum attainable two-sided -value is , limiting the test’s power to detect significance at . Because two pairwise comparisons are reported (REEF vs. NVFlare, REEF vs. LASER), a Bonferroni-adjusted threshold of per comparison is noted; the REEF vs. NVFlare result () remains significant after correction.
Negative control experiments. To validate absence of data leakage and confirm genuine biological signal, we performed negative control experiments where stress labels were randomly permuted before training. The negative control constitutes a true end-to-end permutation test: labels are permuted before any computation begins, encoder warmup training runs against permuted labels, saliency scores are derived from permuted-label-trained encoder weights, feature selection operates on those permuted-label saliency rankings, and federated training proceeds on the resulting selected features — no component of the pipeline has access to true label ordering. Performance on permuted labels provides a baseline for comparison against true-label results. Critically, saliency computation is performed independently within each LOOCV fold using only the 11 training samples for that fold; a fresh encoder is instantiated per fold, warmup training uses only fold-local training indices, and the held-out test sample is excluded before any importance scores are derived, ensuring no test-set information enters the feature selection stage at any point.
Saliency ranking stability analysis. Feature ranking stability was evaluated by computing Jaccard similarity between top-20 feature sets selected at different K thresholds () from a single importance ranking computed on a fixed training set. This analysis tests stability with respect to the selection budget parameter K; because importance scores are computed once from the same ranking, Jaccard=1.0 is mathematically guaranteed when all K values exceed the top-20 threshold, and does not reflect cross-seed or cross-fold stability. Cross-seed and cross-fold stability of selected features represents an important direction for future validation.
@@
All code implemented in Python 3.10 using PyTorch 2.9.1 [18] (neural networks and gradient-based feature importance), NVIDIA FLARE 2.7.1 [15] (VFL infrastructure), scikit-learn 1.7.2 [19] (preprocessing, metrics), NumPy 2.2, and pandas 2.3. Experiments were conducted on consumer-grade hardware (11th Gen Intel Core i7-11800H @ 4.60 GHz, 16 GB RAM, Debian Linux 13, kernel 6.12.57) to demonstrate computational accessibility without specialized GPU requirements. Complete codebase, preprocessed data, and configuration files are available at https://github.com/samvictordr/REEF.